OpenStudy (anonymous):

I need serious help with this question. Find the slope of the line that passes through the origin and (8, -5)

OpenStudy (anonymous):

slope = -5/8

OpenStudy (anonymous):

wow howd you get that?

OpenStudy (anonymous):

well the slope is the gradient of the line. therefore to find the gradient you need to find the difference in y divided by the difference in x. that means that you have to do this: gradient= (Y2 -Y1)/(X2 - X1)...... the origin lies on the points (0,0) , therefore your gradient will be (-5 -0) / (8 -0) = -5/8...